OpenStudy (anonymous):

solve m = 2y + 3y for y. show work please.

OpenStudy (anonymous):

first off, what is \(2y+3y\) ?

OpenStudy (anonymous):

a literal equation...?

OpenStudy (anonymous):

I'm sorry, I stink at math.

OpenStudy (anonymous):

no i mean "what do you get when you add \(2y+3y\) ?

OpenStudy (anonymous):

5y?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so the first step is to write \[m=5y\]

OpenStudy (anonymous):

then, since you want \(y\) and not \(5y\) you want to divide both sides by \(5\)

OpenStudy (anonymous):

ok.. :)

OpenStudy (anonymous):

when i say "divide" i don't mean literally divide, i mean just write it:\[m=5y\] \[\frac{m}{5}=\frac{\cancel{5}y}{\cancel{5}}\] \[\frac{m}{5}=y\]

OpenStudy (anonymous):

wow thanks! that's the completed answer?

OpenStudy (anonymous):

yes, when it is says "solve for \(y\) " that means your final answer should look like \[y=\text{something}\]

OpenStudy (anonymous):

in this case it is \[y=\frac{m}{5}\]